OpenStudy (anonymous):

solve for the variable. reduce the answer to lowest terms 8b=32

OpenStudy (anonymous):

b = 4

jhonyy9 (jhonyy9):

divide both sides by 8

OpenStudy (anonymous):

b=32/8 b = 4

OpenStudy (anonymous):

divide both side by 8 to get b by itself 32/8= 4 check: 4*8=32